Four coins are tossed. Let Z be the random variable representing the number heads that occur. Find the values of the random variable Z. ( Take note that you should have 2 = 16 possible outcomes ).


1
Expert's answer
2022-03-01T10:42:19-0500

There are "2^4=16" possible outcomes


"S=\{HHHH, HHHT, HHTH, HTHH, THHH,"

"HHTT, HTHT, HTTH, THTH, THHT, TTHH,"

"HTTT, THTT, TTHT, TTTH, TTTT\}"

The possible values of the random variable "Z" are "0, 1, 2, 3, 4."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2""\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & Z \\ \hline
 HHHH & 4 \\
 \hdashline
 HHHT & 3 \\
 \hdashline
 HHTH & 3 \\
 \hdashline
 HTHH & 3 \\
 \hdashline
 THHH & 3 \\
 \hdashline
 HHTT & 2 \\
 \hdashline
 HTHT & 2 \\
 \hdashline
 HTTH & 2 \\
 \hdashline
 THTH & 2 \\
 \hdashline
 THHT & 2 \\
 \hdashline
 TTHH & 2 \\
 \hdashline
 HTTT & 1 \\
 \hdashline
 THTT & 1 \\
 \hdashline
 TTHT & 1 \\
 \hdashline
 TTTH & 1 \\
 \hdashline
 TTTT & 0 \\
 \hdashline
\end{array}"


Construct the probability distribution of the random variable


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 z & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(z) & 1/16 & 1/4 & 3/8 & 1/4 & 1/16
\end{array}"

or


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 z & 0 & 1 & 2 & 3 & 4 \\ \hline
 p(z) & 0.0625 & 0.25 & 0.375 & 0.25 & 0.0625
\end{array}"